OpenStudy (braydonlevi99):

Find the coordinates of the midpoint of the segment whose endpoints are H(6,4) and K(2,8). All I need is to know how to do it and know what to do.

pooja195 (pooja195):

Use the midpoint formula \[\huge~\rm~\frac{ x_2+x_1 }{ 2},\frac{ y_2+y_1 }{ 2 } \]

OpenStudy (braydonlevi99):

Okay, thank you so much!

pooja195 (pooja195):

yw ^_^

OpenStudy (braydonlevi99):

\[8+8\div2 = 16\div2 = 8\]

OpenStudy (braydonlevi99):

\[8+2\div2 = 10\div2 = 5\]

OpenStudy (braydonlevi99):

So isn't the answer (8,5)

OpenStudy (braydonlevi99):

My bad I did the wrong one!!

pooja195 (pooja195):

Nope \[\huge~\rm~\frac{ 6+2}{ 2},\frac{ 4+8 }{ 2 }\]

OpenStudy (braydonlevi99):

I did the next one on the study guide which says, "Find the distance between points P(8,) and Q(3,8) to the nearest tenth. Which on that one I would have to use the distance formula.

pooja195 (pooja195):

Lol i see :P Ok so just look at what i gave you above and solve

OpenStudy (braydonlevi99):

The answer is (4,6)

pooja195 (pooja195):

Well done =)

OpenStudy (braydonlevi99):

Mind if I fan you?

pooja195 (pooja195):

lol not at all go for it :P